# Inequality Restricted Minimum Density Power Divergence Estimation in Panel Count Data

### v1

@@ 3 Minimum Density Power Divergence Estimator
In a study of robust estimation, Basu et al. 1998 were the first to propose the density power divergence method. They considered the parametric model of densities $\{f_{t}\}$ with respect to the Lebesgue measure where the unknown parameter $t\in\Theta$ where $\Theta$, the parameter space. Let $G$ be the class of all distributions with densities $g$ for the same measure. Under this assumption, the density power divergence measure between density functions $g$ and $f_{t}$ is defined as,
[…]
The hyperparameter, $\gamma,$ termed as a tuning parameter, plays an essential role in estimating the unknown model parameters. It governs the trade-off between model efficiency and robustness in estimation based on the density power divergence measure. Further, $d_{\gamma}(g,f_{t})$ tends to become the Kullback-Leibler divergence between $g$ and $f_{t}$ when $\gamma$ tends to 0. For estimation purposes, $f$ is set as the assumed theoretical model where true density $g$ is replaced by the empirical one, say $\widehat{g}$. In the current model set-up, the density power divergence between the theoretical probability mass function $f$ and the empirical probability mass function $\widehat{g}$ can be found as follows,
$d_{\gamma}(\widehat{g},f)$ $=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{1+\gamma}(\bm{n})-\left(1+\frac{1}{\gamma}\right)\frac{1}{m}\sum_{i=1}^{m}f^{\gamma}(\bm{N_{i}}),\quad\gamma>0$ (3.2)
where $\bm{n}=(n_{11},n_{21},\ldots,n_{1k},n_{2k}).$ For a given tuning parameter ${\gamma}$, minimizing $d_{\gamma}(\widehat{g},f)$ with respect to the parameter ${\bm{\theta}}$ is the same as minimizing $\frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}})$ where,
$V_{\bm{\theta}}(\bm{N_{i}})=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{1+\gamma}(\bm{n})-\left(1+\frac{1}{\gamma}\right)f^{\gamma}(\bm{N_{i}})$ (3.3)
@@
$-\frac{1}{m}\sum_{i=1}^{m}\log f(\bm{n});$
the relevant minimizer is identified as the maximum likelihood estimator (MLE) of ${\bm{\theta}}$. Based on (3.2), and (3.3), the minimum density power divergence estimator of $\bm{\theta}$ can be derived as,
${\hat{\bm{\theta}}}_{\gamma}=\underset{\bm{\theta}}{\arg\min}\ \frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}});\qquad\gamma>0$ (3.4)
@@ 4 Restricted Minimum Density Power Divergence Estimator
$\bm{h}\left(\bm{\theta}\right)\geqq\mathbf{0}_{r}$ (4.1)
@@
$\mathbf{H}(\bm{\theta})=\frac{\partial\bm{h}^{T}(\bm{\theta})}{\partial\bm{\theta}}$ (4.2)
exists, is continuous in $\bm{\theta}$, and satisfies $\operatorname{rank}(\mathbf{H}(\bm{\theta}))=r$. Under the model assumption above, the restricted MDPD (RMDPD) estimator can be derived as,
${\tilde{\bm{\theta}}}_{\gamma}=\underset{\bm{h(\theta)\geq 0_{r}}}{\arg\min}\ \frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}});\qquad\gamma>0.$ (4.3)
Further, leveraging the application of the Karush–Kuhn–Tucker (KKT) conditions, any value of $\bm{\theta}$ satisfying the following system of equations and inequalities will be the minimizer of (3.3) with the restriction $\bm{h}\left(\bm{\theta}\right)\geq\mathbf{0}_{r}.$
$\frac{\partial}{\partial\bm{\theta}}\frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}})+\bm{\lambda}\bm{H}\left(\bm{\theta}\right)=\mathbf{0}_{p}\$ (4.4)
$\bm{h}\left(\bm{\theta}\right)\geqq\mathbf{0}_{r}\$ (4.5)
$\bm{\lambda}^{\prime}\bm{h}\left(\bm{\theta}\right)=\mathbf{0}\$ (4.6)
$\bm{\lambda}\geqq\mathbf{0}\$ (4.7)
@@
$\bm{A}\bm{\theta}\geqq\mathbf{0}_{r}\$
where,
$\bm{A}=\left(\begin{array}[]{ccccc}0&1&0&-1&0\\ 0&0&1&0&-1\end{array}\right)\$
@@ 4.1. Asymptotic Properties
The RMDPD estimator cannot be derived in explicit form, making the examination of its exact sample properties quite challenging. This leads us to investigate the asymptotic properties of RMDPDE. For further development, we modified the regulatory conditions suggested by Basu et al. 2018 according to our context.
(R1) For every $\bm{\theta}\in\Theta,$ the model distributions $f(\bm{n})$ have a common support, ensuring that the set $\mathcal{S}=\left\{\bm{n}\mid f(\bm{n})>0\right\}$ remains independent of $\bm{\theta}$. Additionally, the true distribution $G$ is supported within $\mathcal{S}$, where the corresponding density ${g}$ remains strictly positive.
(R2) There exists an open subset $\omega$ of the parameter space $\Theta$, consisting of the best fitting parameter $\bm{\theta}_{0}$ such that for nearly all $\bm{n}\in\mathcal{S}$ and every $\bm{\theta}\in\omega$, the density $f(\bm{n})$ is three times differentiable with respect to $\bm{\theta}$, and its third-order partial derivatives are continuous in terms of $\bm{\theta}$.
(R3) The summations $\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{\gamma+1}(\bm{n})$ and $\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{\gamma}(\bm{n})g(\bm{n})$ can be differentiated three times with respect to $\bm{\theta}$, where $g$ represents the true joint mass function of $\bm{N}$. Moreover, these derivatives can be taken directly under the summation sign.
(R4) The $p\times p$ matrix $\bm{J}_{\gamma}(\bm{\theta})$, as specified in equation (4.6), is positive definite.
(R5) There is a function $M_{tuv}(\bm{n})$ such that $\left|\nabla_{tuv}V_{\bm{\theta}}(\bm{n})\right|\leq M_{tuv}(\bm{n})$ for all $\bm{\theta}\in\omega$, where $E_{g}\left[M_{tuv}(\bm{N})\right]=m_{tuv}<\infty$ for all $t,u$, and $v.$
@@
a) The set of equations (4.4) - (4.7) for estimating restricted minimum density power divergence exhibit a consistent sequence of roots $\tilde{\bm{\theta}}_{m,\gamma}$ (hereafter referred to as $\tilde{\bm{\theta}}_{\gamma}$ ), that is, $\tilde{\bm{\theta}}_{\gamma}\xrightarrow[m\rightarrow\infty]{\mathcal{P}}\bm{\theta}_{0}$.
b) The asymptotic distribution of the restricted minimum density power divergence estimator $\tilde{\bm{\theta}}_{\gamma}$ is expressed as
$\sqrt{m}\left(\tilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)\sim N(\mathbf{0}_{p},\bm{\Sigma}_{\gamma}(\bm{\theta_{0}}))$
where, $\bm{\Sigma}_{\gamma}(\bm{\theta_{0}})=L(\bm{\theta_{0}})K_{\gamma}(\bm{\theta_{0}})L(\bm{\theta_{0}})$,
$\bm{L}(\bm{\theta_{0}})=-[J^{-1}_{\gamma}(\bm{\theta_{0}})+J^{-1}_{\gamma}(\bm{\theta_{0}})(H_{s}(\bm{\theta_{0}})P^{-1}H^{T}_{s}(\bm{\theta_{0}})+US^{-1}V)J^{-1}_{\gamma}(\bm{\theta_{0}})]$,
$\bm{J}_{\gamma}(\bm{\theta})=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})\bm{u}^{T}(\bm{n})f^{\gamma+1}(\bm{n}),$ (4.8)
$\bm{K}_{\gamma}(\bm{\theta})=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})\bm{u}^{T}(\bm{n})f^{2\gamma+1}(\bm{n})-\bm{\xi}_{\gamma}(\bm{\theta})\bm{\xi}_{\gamma}^{T}(\bm{\theta}),$ (4.9)
@@ 5 Optimal tuning Parameter
In estimation based on the density power divergence measure, the tuning parameter $\gamma$ is crucial for balancing the trade-off between efficiency and robustness of the RMDPD estimator in the presence of data contamination. Selecting an optimal value for $\gamma$ is vital to ensure the robustness of the estimator without compromising its efficiency. In the literature, several studies have been performed in this direction. Warwick and Jones 2005 proposed a data-driven procedure to achieve the optimal tuning parameter for the MDPDE by minimizing its asymptotic mean squared error (MSE). The expression for this minimization is defined as,
@@
where, $\bm{\theta_{P}}$ denotes the pilot estimator, and $\operatorname{Tr}(\cdot)$ represents the trace of the matrix. However, the procedure introduced by Warwick and Jones 2005 has a limitation, as the optimal value of $\gamma$ is dependent on the choice of the pilot estimator. To mitigate this issue, Basak et al. 2021 introduced an iterative approach. This is referred to as the Iterative Warwick and Jones approach, more commonly known as the ”IWJ” technique. In this method, the pilot estimator is substituted at every iteration with the revised estimator calculated using the optimal value of $\gamma$, and this process continues until stabilization is achieved.
@@
Recently, Sugasawa and Yonekura 2021 delved into this problem by exploiting the asymptotic approximation of Hyvärinen score with unnormalized models based on robust divergence. This criterion eliminates the need for the estimator’s asymptotic variance formula, which is required in the previous method. The Hyvärinen score-matching method, proposed by Hyvärinen and Dayan 2005, is designed to estimate statistical models with an unknown normalization constant. Rather than relying on traditional approaches such as Markov Chain Monte Carlo (MCMC) or approximating the normalization constant, this method minimizes the expected squared distance between the gradient of the log density derived from the model and that obtained from the observed data. However, this method is only applicable when all variables have a continuous value and are defined over $\mathbb{R}^{n}$. Consequently, Hyvärinen 2007 expanded this framework to accommodate binary variables and to estimate non-normalized models defined within the non-negative real domain, denoted as +$\mathbb{R}^{n}_{+}$. Later, Lyu 2012 introduced a generalized version of score matching by incorporating general linear operators into the Fisher divergence. He also presented a specific application of this generalized score-matching to discrete data, demonstrating it as a more intuitive extension of score-matching for discrete cases. Recently, Xu et al. 2022 introduced a generalized score-matching method tailored for regression-type models with count data. They also developed a comprehensive theoretical framework for estimating score matching in cases involving independent but not necessarily identically distributed (INID) observations within a fixed design set-up for regression-type models. This method is referred to as the ”GSM” method, for short.
@@
In this work, our novel contribution includes determining the optimal value of the tuning parameter $\gamma$ through a generalized score-matching technique, proposed by Xu et al. 2022. The following developments are made in this direction. As discussed in 2, in this study we get $m$ independent observation vectors $(n_{i11},n_{i21},\ldots,n_{i1k},n_{i2k})$ for $i=1,\ldots,m.$
Define, $p(\bm{n_{i}}|\bm{\theta})$ as
@@
$p(\bm{n_{i}}|\bm{\theta})=\exp(-(\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{1+\gamma}(\bm{n})-\left(1+\frac{1}{\gamma}\right)f^{\gamma}(\bm{n_{i}})))$ (5.1)
@@
Note that, $p(\bm{n}|\bm{\theta})$ poses to be a parametrized model without the normalization constant. Assume $q(\bm{n})$ be the true probability model, where $q(\bm{n})\propto p(\bm{n})$, when the parametrized model is correctly specified. According to Lyu 2012, a generalized form of score matching can be achieved by substituting the gradient operator with a general linear operator $\mathcal{L}$. As a result, the generalized score matching objective function, denoted by $D_{\mathcal{L}}(q,p)$, is defined as follows,
$D_{\mathcal{L}}(q,p)=\mathbb{E}\left[\left\|\frac{\mathcal{L}q(\bm{n})}{q(\bm{n})}-\frac{\mathcal{L}p(\bm{n}|\bm{\theta})}{p(\bm{n}|\bm{\theta})}\right\|^{2}\right],$ (5.2)
[…]
$\mathcal{L}p\left(\bm{n}\mid\bm{\theta}\right)=\left(\begin{array}[]{c}\vdots\\ \mathcal{L}_{s}p\left(\bm{n}\mid\bm{\theta}\right)\\ \vdots\end{array}\right)=\left(\begin{array}[]{c}\vdots\\ p\left(\bm{n}^{(s+)}\mid\bm{\theta}\right)-p\left(\bm{n}\mid\bm{\theta}\right)\\ \vdots\end{array}\right).$
@@
and $\bm{n}^{(s+)}$ is derived from $\bm{n}$ by adding 1 to its $s^{th}$ element without altering the other elements. Similarly, $\bm{n}^{(s-)}$ can be obtained by subtracting 1 from its $s^{th}$ element while leaving the other elements unchanged. If $\bm{n}$ reaches the boundary of its range, we set $p\left(\bm{n}^{(s+)}\mid\bm{\theta}\right)=0$, and $p\left(\bm{n}^{(s-)}\mid\bm{\theta}\right)=0$. In the discrete multivariate scenario, the operator $\mathcal{L}p\left(\bm{n}\mid\bm{\theta}\right)$ serves as a discrete approximation to the gradient of $p\left(\bm{n}\mid\bm{\theta}\right)$ at $\bm{n}$. By rewriting
@@
In order to handle potential division by zero in these slope expressions, we adopt the following transformation proposed by Hyvärinen 2007,

### v4

@@ 3 Minimum Density Power Divergence Estimator
In a study of robust estimation, Basu et al. [19] were the first to propose the density power divergence method. They considered the parametric model of densities $\{f_{t}\}$ with respect to the Lebesgue measure where the unknown parameter $t\in\Theta$ where $\Theta$, the parameter space. Let $G$ be the class of all distributions with densities $g$ for the same measure. Under this assumption, the density power divergence measure between density functions $g$ and $f_{t}$ is defined as
$d_{\gamma}(g,f_{t})=\int\left\{f_{t}^{1+\gamma}(x)-\left(1+\frac{1}{\gamma}\right)g(x)\,f_{t}^{\gamma}(x)+\frac{1}{\gamma}g^{1+\gamma}(x)\right\}dx,\quad\gamma>0.$ (3)
The hyperparameter, $\gamma,$ termed as a tuning parameter, plays an essential role in estimating the unknown model parameters. It governs the trade-off between model efficiency and robustness in estimation based on the density power divergence measure. Further, $d_{\gamma}(g,f_{t})$ tends to become the Kullback-Leibler divergence between $g$ and $f_{t}$ when $\gamma$ tends to 0. For estimation purposes, $f$ is set as the assumed theoretical model where true density $g$ is replaced by the empirical one, say $\widehat{g}$. In the current model set-up, the density power divergence between the theoretical probability mass function $f$ and the empirical probability mass function $\widehat{g}$ can be found as follows
$d_{\gamma}(\widehat{g},f)$ $=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{1+\gamma}(\bm{n})-\left(1+\frac{1}{\gamma}\right)\frac{1}{m}\sum_{i=1}^{m}f^{\gamma}(\bm{N_{i}}),\quad\gamma>0$ (4)
where $\bm{n}=(n_{11},n_{21},\ldots,n_{1k},n_{2k}).$ For a given tuning parameter $\gamma$}, minimizing $d_{\gamma}(\widehat{g},f)$ with respect to the parameter $\bm{\theta}$} is the same as minimizing $\frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}})$ where
$V_{\bm{\theta}}(\bm{N_{i}})=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{1+\gamma}(\bm{n})-\left(1+\frac{1}{\gamma}\right)f^{\gamma}(\bm{N_{i}}).$ (5)
@@
$\frac{1}{m}\sum_{i=1}^{m}\log f(\bm{n});$
the relevant minimizer is identified as the maximum likelihood estimator (MLE) of $\bm{\theta}$}. Based on (4), and (5), the minimum density power divergence estimator of $\bm{\theta}$ can be derived as
${\hat{\bm{\theta}}}_{\gamma}=\underset{\bm{\theta}}{\arg\min}\ \frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}}),\qquad\gamma>0.$ (6)
@@ 4 Restricted Minimum Density Power Divergence Estimator
$\bm{h}\left(\bm{\theta}\right)\geq\mathbf{0}_{r}$ (7)
@@
$\mathbf{H}(\bm{\theta})=\frac{\partial\bm{h}^{T}(\bm{\theta})}{\partial\bm{\theta}}$ (8)
exists, is continuous in $\bm{\theta}$, and satisfies $\operatorname{rank}(\mathbf{H}(\bm{\theta}))=r$. Under the model assumption above, the restricted MDPD (RMDPD) estimator can be derived as
${\tilde{\bm{\theta}}}_{\gamma}=\underset{\bm{h(\theta)\geq 0_{r}}}{\arg\min}\ \frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}}),\qquad\gamma>0.$ (9)
Further, leveraging the application of the Karush–Kuhn–Tucker (KKT) conditions, any value of $\bm{\theta}$ satisfying the following system of equations and inequalities will be the minimizer of (5) with the restriction $\bm{h}\left(\bm{\theta}\right)\geq\mathbf{0}_{r},$
$\frac{\partial}{\partial\bm{\theta}}\frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}})+\bm{\lambda}\bm{H}\left(\bm{\theta}\right)=\mathbf{0}_{p},\\$ (10)
$\bm{h}\left(\bm{\theta}\right)\geq\mathbf{0}_{r},\\$ (11)
$\bm{\lambda}^{\prime}\bm{h}\left(\bm{\theta}\right)=\mathbf{0},\\$ (12)
$\bm{\lambda}\geq\mathbf{0},\\$ (13)
@@
$\bm{A}\bm{\theta}\geq\mathbf{0}_{r},$
where
$\bm{A}=\left(\begin{array}[]{ccccc}0&1&0&-1&0\\ 0&0&1&0&-1\end{array}\right).\\$
@@
4.1. Computational Algorithm
The objective function derived in this work may exhibit either convexity or non-convexity. To effectively handle such cases, we exploit the computationally efficient framework of Sequential Convex Programming (SCP), as introduced by Dinh and Diehl [29]. SCP is an iterative optimization methodology designed for non-convex problems, wherein each iteration involves constructing a convex surrogate–typically via linearization or convexification–of the original objective and constraints around the current estimate of the solution. This enables the application of convex optimization techniques to refine the solution progressively. Owing to its robustness and real-time feasibility, SCP has gained widespread use in domains such as trajectory optimization, control systems, and robotics.
SCP is widely known for its capability to tackle intricate, constrained problems where other straightforward non-convex optimization techniques would be too computationally intensive. Here, to find the optimal values of $\bm{\theta}$, our strategy is to consider the classical gradient; see Duchi et al. [40]. Let us assume
$f(\bm{\theta})=\frac{1}{m}\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}}).$
Therefore, the MDPD estimation is reduced to a general non-convex optimization problem as follows,
$\min_{\bm{\theta}\in\mathbb{R}^{p}_{+}}$ $f(\bm{\theta})$ (14)
subject to $h^{*}_{j}(\bm{\theta})\leq 0,\quad j=1,\ldots,r,$
where $h^{*}_{j}(\bm{\theta})=-h_{j}(\bm{\theta})$. At any $k^{th}$ iteration, the SCP approach maintains an estimate of the solution $\bm{\theta}^{(k)}$, and a trust region $\mathcal{T}^{(k)}\subset\mathbb{R}^{p}_{+}$ over which the approximations are deemed valid. The trust region method is a classical tool within SCP and typically involves an $\ell_{2}$-norm ball
$\mathcal{T}^{(k)}=\left\{\bm{\theta}\in\mathbb{R}^{p}\,\big|\,\|\bm{\theta}-\bm{\theta}^{(k)}\|_{2}\leq\rho\right\},\quad\rho>0$
or an $\ell_{1}$-norm box
$\mathcal{T}^{(k)}=\left\{\bm{\theta}\in\mathbb{R}^{p}\,\big|\,|\bm{\theta_{i}}-\bm{\theta_{i}}^{(k)}|\leq\rho_{i},\quad i=1,\ldots\,p\right\}.$
Here, we see that the $\ell_{2}$-norm ball trust region offers greater flexibility in the development of modeling strategies. In contrast, the $\ell_{1}$-norm box trust region limits the indices $\bm{\theta_{i}}$ to engage solely in convex objectives, convex inequality constraints, and unconstrained linear equality constraints (i.e., $\rho_{i}=+\infty$). Thereafter, we form a convex approximation of $f$ applying a first-order Taylor approximation as follows
$\widehat{f}(\bm{\theta})=f(\bm{\theta}^{(k)})+\nabla f(\bm{\theta}^{(k)})^{T}(\bm{\theta}-\bm{\theta}^{(k)}).\\$ (15)
Similarly, we can proceed with the convex approximations $\tilde{h}^{*}_{j}$ of the inequality constraints $h^{*}_{j}$ over $\mathcal{T}^{(k)}$ for $j=1,\dots,r.$
Finally, we can solve the convex subproblem as follows,
$\min_{\bm{\theta}\in\mathcal{T}^{(k)}}$ $\widehat{f}(\bm{\theta})$ (16)
subject to $\tilde{h}^{*}_{j}(\bm{\theta})\leq 0,\quad j=1,\ldots,r.$
The steps for SCP are provided in the following algorithm.
Algorithm: Sequential Convex Programming
Initialize $\bm{\theta}^{(1)}$ and set the trust region radius $\rho_{i}>0$, $i=1,\ldots,p$. Iteration: - Set $k=1$. - Form the affine (first-order Taylor) approximation around $\bm{\theta}^{(k)}$: $\widehat{f}(\bm{\theta})=f(\bm{\theta}^{(k)})+\nabla f(\bm{\theta}^{(k)})^{T}(\bm{\theta}-\bm{\theta}^{(k)}).$ - Define the trust region: $\mathcal{T}^{(k)}=\left\{\bm{\theta}\in\mathbb{R}^{p}\,\big|\,|\bm{\theta}-\bm{\theta}^{(k)}|\leq\rho_{i},\quad i=1,\ldots\,p\right\}.$ - Construct convex approximations $\tilde{h}^{*}_{j}(\bm{\theta})$ of the constraints $h^{*}_{j}(\bm{\theta})$, for $j=1,\ldots,r$. - Solve the convex subproblem: $\min_{\bm{\theta}\in\mathcal{T}^{(k)}}$ $\widehat{f}(\bm{\theta})$ subject to $\tilde{h}^{*}_{j}(\bm{\theta})\leq 0,\qquad j=1,\ldots,r.$ Update: - Set $k=k+1$ and $\bm{\theta}^{(k+1)}$ as the optimal solution of the above subproblem. Repeat: Until convergence Return $\bm{\theta}^{(k+1)}$ as the optimal estimate.
@@ 4.2. Asymptotic Properties
The RMDPD estimator cannot be derived in explicit form, making the examination of its exact sample properties quite challenging. This leads us to investigate the asymptotic properties of RMDPDE. For further development, we modified the regulatory conditions suggested by Basu et al. [26] according to our context.
  • (R1)
    For every $\bm{\theta}\in\Theta,$ the model distributions $f(\bm{n})$ have a common support, ensuring that the set $\mathcal{S}=\left\{\bm{n}\mid f(\bm{n})>0\right\}$ remains independent of $\bm{\theta}$. Additionally, the true distribution $G$ is supported within $\mathcal{S}$, where the corresponding density g{g} remains strictly positive.
  • (R2)
    There exists an open subset $\omega$ of the parameter space $\Theta$, consisting of the best fitting parameter $\bm{\theta}_{0}$ such that for nearly all $\bm{n}\in\mathcal{S}$ and every $\bm{\theta}\in\omega$, the density $f(\bm{n})$ is three times differentiable with respect to $\bm{\theta}$, and its third-order partial derivatives are continuous in terms of $\bm{\theta}$.
  • (R3)
    The summations $\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{\gamma+1}(\bm{n})$ and $\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{\gamma}(\bm{n})g(\bm{n})$ can be differentiated three times with respect to $\bm{\theta}$, where $g$ represents the true joint mass function of $\bm{N}$. Moreover, these derivatives can be taken directly under the summation sign.
  • (R4)
    The $p\times p$ matrix $\bm{J}_{\gamma}(\bm{\theta})$, as specified in equation (12), is positive definite.
  • (R5)
    There is a function $M_{tuv}(\bm{n})$ such that $\left|\nabla_{tuv}V_{\bm{\theta}}(\bm{n})\right|\leq M_{tuv}(\bm{n})$ for all $\bm{\theta}\in\omega$, where $E_{g}\left[M_{tuv}(\bm{N})\right]=m_{tuv}<\infty$ for all $t,u$, and $v.$
@@
  • a)
    The set of equations (10) - (13) for estimating restricted minimum density power divergence exhibit a consistent sequence of roots $\tilde{\bm{\theta}}_{m,\gamma}$ (hereafter referred to as $\tilde{\bm{\theta}}_{\gamma}$ ), that is, $\tilde{\bm{\theta}}_{\gamma}\xrightarrow[m\rightarrow\infty]{\mathcal{P}}\bm{\theta}_{0}$.
  • b)
    The asymptotic distribution of the restricted minimum density power divergence estimator $\tilde{\bm{\theta}}_{\gamma}$ is expressed as
    $\sqrt{m}\left(\tilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)\sim N(\mathbf{0}_{p},\bm{\Sigma}_{\gamma}(\bm{\theta_{0}})),$
    where $\bm{\Sigma}_{\gamma}(\bm{\theta_{0}})=L^{T}(\bm{\theta_{0}})K_{\gamma}(\bm{\theta_{0}})L(\bm{\theta_{0}})$,
    $\bm{L}(\bm{\theta_{0}})=-[J^{-1}_{\gamma}(\bm{\theta_{0}})+J^{-1}_{\gamma}(\bm{\theta_{0}})(H_{s}(\bm{\theta_{0}})P^{-1}H^{T}_{s}(\bm{\theta_{0}}))J^{-1}_{\gamma}(\bm{\theta_{0}})]$,
    $\bm{J}_{\gamma}(\bm{\theta})=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})\bm{u}^{T}(\bm{n})f^{\gamma+1}(\bm{n}),$ (17)
    $\bm{K}_{\gamma}(\bm{\theta})=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})\bm{u}^{T}(\bm{n})f^{2\gamma+1}(\bm{n})-\bm{\xi}_{\gamma}(\bm{\theta})\bm{\xi}_{\gamma}^{T}(\bm{\theta}),$ (18)
@@ 5 Optimal tuning Parameter
In estimation based on the density power divergence measure, the tuning parameter $\gamma$ is crucial for balancing the trade-off between efficiency and robustness of the RMDPD estimator in the presence of data contamination. Selecting an optimal value for $\gamma$ is vital to ensure the robustness of the estimator without compromising its efficiency. Several studies have been performed in the literature in this direction. Warwick and Jones [33] proposed a data-driven procedure to achieve the optimal tuning parameter for the MDPDE by minimizing its asymptotic mean squared error (MSE). The expression for this minimization is defined as
@@
where $\bm{\theta_{P}}$ denotes the pilot estimator, and $\operatorname{Tr}(\cdot)$ represents the trace of the matrix. However, the procedure introduced by Warwick and Jones [33] has a limitation, as the optimal value of $\gamma$ is dependent on the choice of the pilot estimator. To mitigate this issue, Basak et al. [41] introduced an iterative approach. This is referred to as the Iterative Warwick and Jones approach, more commonly known as the ”IWJ” technique. In this method, the pilot estimator is substituted at every iteration with the revised estimator calculated using the optimal value of $\gamma$, and this process continues until stabilization is achieved.
@@
Recently, Sugasawa and Yonekura [34] delved into this problem by exploiting the asymptotic approximation of Hyvärinen score with unnormalized models based on robust divergence. This criterion eliminates the need for the estimator’s asymptotic variance formula, which is required in the previous method. The Hyvärinen score-matching method, proposed by Hyvärinen and Dayan [36], is designed to estimate statistical models with an unknown normalization constant. Rather than relying on traditional approaches such as Markov Chain Monte Carlo (MCMC) or approximating the normalization constant, this method minimizes the expected squared distance between the gradient of the log density derived from the model and that obtained from the observed data. However, this method is only applicable when all variables have a continuous value and are defined over $\mathbb{R}^{n}$. Consequently, Hyvärinen [37] expanded this framework to accommodate binary variables and to estimate non-normalized models defined within the non-negative real domain, denoted as $\mathbb{R}^{n}_{+}$. Later, Lyu [38] introduced a generalized version of score matching by incorporating general linear operators into the Fisher divergence. He also presented a specific application of this generalized score-matching to discrete data, demonstrating it as a more intuitive extension of score-matching for discrete cases. Recently, Xu et al. [39] introduced a generalized score-matching method tailored for regression-type models with count data. They also developed a comprehensive theoretical framework for estimating score matching in cases involving independent observations within a fixed design set-up for regression-type models. This method is referred to as the ”GSM” method, for short.
@@
In this work, our novel contribution includes determining the optimal value of the tuning parameter $\gamma$ through a generalized score-matching technique, proposed by Xu et al. [39]. The following developments are made in this direction. As discussed in 2, in this study we get $m$ independent observation vectors $(n_{i11},n_{i21},\ldots,n_{i1k},n_{i2k})$ for $i=1,\ldots,m.$ Define, $p(\bm{n_{i}}|\bm{\theta})$ as
@@
$p(\bm{n_{i}}|\bm{\theta})=\exp\left(-\left(\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}f^{1+\gamma}(\bm{n})-\left(1+\frac{1}{\gamma}\right)f^{\gamma}(\bm{n_{i}})\right)\right)$ (19)
@@
Note that, $p(\bm{n}|\bm{\theta})$ poses to be a parametrized model without the normalization constant. Assume $q(\bm{n})$ be the true probability model, where $q(\bm{n})\propto p(\bm{n})$, when the parametrized model is correctly specified. According to Lyu [38], a generalized form of score matching can be achieved by substituting the gradient operator with a general linear operator $\mathcal{L}$. As a result, the generalized score matching objective function, denoted by $D_{\mathcal{L}}(q,p)$, is defined as follows
$D_{\mathcal{L}}(q,p)=\mathbb{E}\left[\left\|\frac{\mathcal{L}q(\bm{n})}{q(\bm{n})}-\frac{\mathcal{L}p(\bm{n}|\bm{\theta})}{p(\bm{n}|\bm{\theta})}\right\|^{2}\right],$ (20)
where $\|\cdot\|$ denotes the Euclidean norm. Following Xu et al. [39], we define a linear operator $\mathcal{L}$ as
$\mathcal{L}p\left(\bm{n}\mid\bm{\theta}\right)=\left(\begin{array}[]{c}\vdots\\ \mathcal{L}_{s}p\left(\bm{n}\mid\bm{\theta}\right)\\ \vdots\end{array}\right)=\left(\begin{array}[]{c}\vdots\\ p\left(\bm{n}^{(s+)}\mid\bm{\theta}\right)-p\left(\bm{n}\mid\bm{\theta}\right)\\ \vdots\end{array}\right),$
@@
where $\bm{n}^{(s+)}$ is derived from $\bm{n}$ by adding 1 to its $s^{th}$ element without altering the other elements. Similarly, $\bm{n}^{(s-)}$ can be obtained by subtracting 1 from its $s^{th}$ element while leaving the other elements unchanged. If $\bm{n}$ reaches the boundary of its range, we set $p\left(\bm{n}^{(s+)}\mid\bm{\theta}\right)=0$, and $p\left(\bm{n}^{(s-)}\mid\bm{\theta}\right)=0$. In the discrete multivariate scenario, the operator $\mathcal{L}p\left(\bm{n}\mid\bm{\theta}\right)$ serves as a discrete approximation to the gradient of $p\left(\bm{n}\mid\bm{\theta}\right)$ at $\bm{n}$. By rewriting
@@
In order to handle potential division by zero in these slope expressions, we adopt the following transformation proposed by Hyvärinen [37],